# Can Current Detectors Catch Face-to-Voice Deepfake Attacks?

### v1

@@ Appendix A Detectors Accuracy on FOICE Data
Figure 5 shows the classification accuracy of the four detectors across FOICE originals, noisy, and denoised audio conditions. The results demonstrate that FOICE fine-tuning consistently improves accuracy, with AASIST and Ren et al. [15] achieving near-perfect performance, while Sun et al. [18] and TCM also benefit though remain less robust overall.
Figure 5: Accuracy of the four detectors on FOICE originals, noisy, and denoised audio. Solid bars show baseline performance, and hatched bars show fine-tuned variants. After fine-tuning, AASIST and Ren et al. [15] achieve near-perfect accuracy across all conditions, while Sun et al. [18] and TCM improve substantially, though remain weaker overall. The results highlight the effectiveness of fine-tuning in boosting in-distribution detection, even under degraded audio conditions.
@@
The spectrograms in Figure 6 present representative samples from the vocoders contained in the datasets used to train the baseline detectors, with each vocoder exhibiting unique spectral envelopes, excitation characteristics, and artifact structures that are visually distinct in both frequency and time. The training provenance for each detector is as follows: AASIST and TCM were both trained using ASVspoof2019; Sun et al. [18] was trained on LibriSeVoc, ASVspoof2019, and WaveFake; and, following the protocol in [15], Ren et al. [15] was trained in a leave-one-dataset-out manner on LibriSeVoc, ASVspoof2019, WaveFake, and FakeAVCeleb—we utilize the model version trained with LibriSeVoc and ASVspoof2019. The datasets themselves span a broad set of vocoders: LibriSeVoc incorporates DiffWave, MelGAN, Parallel WaveGAN, WaveGrad, WaveNet, WaveRNN; WaveFake includes MelGAN, Parallel WaveGAN, WaveGlow, MultiBand-MelGAN, and FullBand-MelGAN, HiFi-GAN; while ASVspoof2019 contains WaveNet, WaveRNN, Spectral, WORLD, Waveform, Griffin-Lim, MFCC-vocoder, Neural source-filter, STRAIGHT, and VocAine; for FakeAVCeleb, there is only SV2TTS.
Figure 6: Spectrograms of vocoders present in LibriSeVoc, ASVspoof2019, and WaveFake, corresponding to the training sources of AASIST, Sun et al. [18], Ren et al. [15] (variant trained on LibriSeVoc and ASVspoof2019), and TCM; each vocoder displays unique spectrotemporal patterns that baseline detectors may overfit to or miss, depending on dataset composition.

### v2

@@ Appendix A Detectors Accuracy on FOICE Data
Figure 5 shows the classification accuracy of the four detectors across FOICE originals, noisy, and denoised audio conditions. The results demonstrate that FOICE fine-tuning consistently improves accuracy, with AASIST and Ren et al. [AISVGeneralization] achieving near-perfect performance, while Sun et al. [VocoderArtifacts] and TCM also benefit though remain less robust overall.
Figure 5: Accuracy of the four detectors on FOICE originals, noisy, and denoised audio. Solid bars show baseline performance, and hatched bars show fine-tuned variants. After fine-tuning, AASIST and Ren et al. [AISVGeneralization] achieve near-perfect accuracy across all conditions, while Sun et al. [VocoderArtifacts] and TCM improve substantially, though remain weaker overall. The results highlight the effectiveness of fine-tuning in boosting in-distribution detection, even under degraded audio conditions.
@@
The spectrograms in Figure 6 present representative samples from the vocoders contained in the datasets used to train the baseline detectors, with each vocoder exhibiting unique spectral envelopes, excitation characteristics, and artifact structures that are visually distinct in both frequency and time. The training provenance for each detector is as follows: AASIST and TCM were both trained using ASVspoof2019; Sun et al. [VocoderArtifacts] was trained on LibriSeVoc, ASVspoof2019, and WaveFake; and, following the protocol in [AISVGeneralization], Ren et al. [AISVGeneralization] was trained in a leave-one-dataset-out manner on LibriSeVoc, ASVspoof2019, WaveFake, and FakeAVCeleb—we utilize the model version trained with LibriSeVoc and ASVspoof2019. The datasets themselves span a broad set of vocoders: LibriSeVoc incorporates DiffWave, MelGAN, Parallel WaveGAN, WaveGrad, WaveNet, WaveRNN; WaveFake includes MelGAN, Parallel WaveGAN, WaveGlow, MultiBand-MelGAN, and FullBand-MelGAN, HiFi-GAN; while ASVspoof2019 contains WaveNet, WaveRNN, Spectral, WORLD, Waveform, Griffin-Lim, MFCC-vocoder, Neural source-filter, STRAIGHT, and VocAine; for FakeAVCeleb, there is only SV2TTS.
Figure 6: Spectrograms of vocoders present in LibriSeVoc, ASVspoof2019, and WaveFake, corresponding to the training sources of AASIST, Sun et al. [VocoderArtifacts], Ren et al. [AISVGeneralization] (variant trained on LibriSeVoc and ASVspoof2019), and TCM; each vocoder displays unique spectrotemporal patterns that baseline detectors may overfit to or miss, depending on dataset composition.